# Bayesian evidence for two peaks in the sound speed in cold dense QCD

### v3

@@ Abstract
I show that in addition to the well-known peak inside massive neutron stars, the sound speed in cold dense QCD matter likely exhibits another peak above neutron star densities before asymptotic to $c_{s}=\sqrt{C_{s}}=\sqrt{1/3}$. Based on the framework reported in arxiv:2408.16738, this approach does not rely on any assumption about the ultra-dense matter not realized in nature. Current multimessenger observation of neutron stars favors the two-peak scenario with a Bayes factor $5.1_{-0.7}^{+0.9}$, where the uncertainties are systematics due to models of neutron star inner cores. This evidence grows to $27_{-8}^{+11}$ if the $2.6M_{\odot}$ component of GW190814 is a neutron star. The trough in $C_{s}$ separating the two peaks is inferred to be below $0.05^{+0.04}_{-0.02}$ (at the $50\%$ level) if $2.6M_{\odot}$ neutron stars exist. The second peak above $1/3$ beyond baryon chemical potential $\mu_{B}=1.6-1.9$ GeV most likely signals non-perturbative effects in cold quark matter, for instance color superconductivity.
@@ I Introduction
The QCD phase diagram at low temperatures and high baryon density remains elusive because the theoretical and experimental tools needed remain rudimentary. Lattice QCD, the only reliable method to calculate non-perturbative phenomena in QCD, is inapplicable as it suffers from the well-known Fermion sign problem  [1, 2, 3] and heavy-ion collisions cannot simultaneously access low temperatures and high densities. Neutron stars (NSs) provide a unique laboratory to study high-density matter at low-temperature and observation of NS mass, radius, and tidal deformability have provided useful constraints on the equation of state (EOS) at baryon number densities $n_{B}\simeq 2-10n_{0}$ where $n_{0}=0.16~{}\text{fm}^{-3}$ (e.g.,[4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]). Above NS densities reliable descriptions in the range $n_{B}\simeq 10-30n_{0}$ are lacking as the strongly interacting cold quark matter (QCM) renders perturbative calculations untractable. Non-perturbative phenomenon in particular color superconductivity [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28] is believed to be active in this regime, although estimates of its strength vary [29, 30, 31, 32, 33, 34].
@@
I now derive a set of stronger conditions that, at the expense of becoming more conservative, are entirely expressible in terms of the EOS at NS and pQCD densities. To begin with, bounds on the magnitude of $C_{s,\mathrm{min}}^{\mathrm{UD}}$ and $C_{s,\mathrm{max}}^{\mathrm{UD}}$ are reported in the preceding letter [66]. A central ingredient is the mean value EOS, the constant sound speed model that directly connects TOV and pQCD points in the baryon number density $n_{B}$ versus baryon chemical potential $\mu_{B}$ plane. It is shown as the dashed black line in fig. 1. Since slopes of $n_{B}(\mu_{B})$ relations are related to the sound speed as
@@
The mean value bound eq. 6 can be used to rewrite eq. 2 as
@@
This is a stronger requirement for NS EOSs since $C_{s,\mathrm{mean}}^{\mathrm{UD}}\geq C_{s,\mathrm{min}}^{\mathrm{UD*}}\geq C_{s,\mathrm{min}}^{\mathrm{UD}}$.
@@
For a given NS model, the requirements eqs. 10, 2 and 3 can be simplified depending on its prediction for $C_{s,\mathrm{mean}}^{\mathrm{UD}}$. When $C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$, the mean value bound implies eq. 3 if $C_{s,\mathrm{min}}^{\mathrm{UD}}=C_{s,\mathrm{min}}^{\mathrm{UD*}}$, so eqs. 10 and 7 alone suffice in this scenario. I prove in appendix A that this statement holds even if $C_{s,\mathrm{max}}^{\mathrm{UD}}$ is not the global maximum $C_{s,\mathrm{max}}^{\mathrm{UD*}}$.
@@
when $C_{s,\mathrm{mean}}^{\mathrm{UD}}<1/3$. A proof is presented in appendix A. Intuitively, eq. 12 assures a second peak because EOSs obeying it are soft in the ultra-dense regime and must have segments resembling the maximally soft model characterized by a low trough preceding a high peak.
@@ III Bayesian evidence
To assess the evidence for eq. 13 I adopt a Bayesian approach where ab-initio nuclear theories are taken as the prior knowledge and astrophysical observations are imposed as constraints. Outer layers of NSs are constructed following the prescription in refs [83, 64]. It centers on ${\chi\text{EFT}}$ calculations of the pure neutron matter up to N3LO  [84, 85] along with $2\sigma$ truncation errors and is informed by empirical constraints on the symmetric nuclear matter. NS inner cores where the density exceeds $n_{\chi\text{EFT}}=2n_{0}$ are modeled by a class of agnostic parameterizations detailed in appendix B. As for the CQM EOS in the perturbative regime, I take N2LO pQCD predictions [51, 50] at $\mu_{B}=2.4$ GeV supplemented by the soft contribution $\sim\mathcal{O}(\alpha_{s}^{3}\log^{2}\alpha_{s})$ at N3LO [56, 57].
@@
in the absence of superconducting gaps [64, 65]. This requirement is the result of causality and thermodynamic stability, and ensures every sample is physical. To address the phase space volume, I impose equally probable distributions on the criterion $\Delta P-\Delta P_{\mathrm{mean,max}}$ that are flat over each side ($+$ and --), yielding neutral priors without preferences. Other choices of priors are discussed in appendix C.
Table 1 shows Bayes evidence for scenarios where two peaks are required. It is sensitive to the pQCD renormalization scale $X=\bar{\Lambda}/(\mu_{B}/3)$ and reveals diametric predilections in the extreme cases $X=1,4$. Larger $X$’s lend stronger support because they predict higher $P_{\text{pQCD}}$ thus higher $\Delta P$. Marginalization over this uncertainty is achieved by sampling X from a uniform distribution over the commonly assumed range $X\in[1,4]$ and more restrictive priors motivated by the fact that $P_{\text{pQCD}}$ near $X=1$ is susceptible to pairing contributions [64], by resummed leading and next-to-leading soft contributions [55, 59], and by hints from lattice calculations at finite isospin [86, 33, 34]. The results generally favor the two-peak scenario, and are predominantly determined by the NS maximum mass $M_{\text{TOV}}$ where Bayes factors grow rapidly with increasing $M_{\text{TOV}}$. Constraints on NS radii have very limited effects because current uncertainties are large, and that NS radii may not robustly inform the EOS near maximum mass except in extreme scenarios [64]. Gravitational wave constraints on tidal deformability $\Lambda_{1.4M_{\odot}}\lesssim 600-800$ are implied by the N3LO ${\chi\text{EFT}}$ EOS up to $2n_{0}$ [49, 64]. Results employing ${\chi\text{EFT}}$ to lower densities are reported in appendix C.
@@
This framework also enables inferring the height of the second peak and the preceding trough. Here, in addition to the physical requirement eq. 14, I impose a similar constraint where $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ is sampled randomly from a uniform distribution on $[0,1]$. This injected likelihood excludes NS EOSs demanding $C_{s}^{\mathrm{UD}}>C_{s,\mathrm{max}}^{\mathrm{UD*}}$, so the resulting posterior on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ provides an estimate for the upper bound on the maximum sound speed  [^2] To be clear, $C_{s}^{\mathrm{UD}}$ may surpass this bound but is not required to.. An upper limit on $C_{s,\mathrm{min}}^{\mathrm{UD*}}$ is then obtained for each NS model by solving $\Delta P=\Delta P_{\mathrm{max}}({\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD*}}\right\}},C_{s,\mathrm{max}}^{\mathrm{UD*}})$ when $\Delta P>\Delta P_{\mathrm{mean}}$ and $\Delta P=\Delta P_{\mathrm{min}}({\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD*}}\right\}},C_{s,\mathrm{max}}^{\mathrm{UD*}})$ otherwise, see ref [66].
Figure 2 shows constraints on the magnitude of the peak and the trough. While posteriors on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ only marginally deviate from the prior suggesting limited constraints on this upper bound, the minimum sound speed is informed. For $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, one finds ${\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD*}}\right\}}=0.05^{+0.08}_{-0.04}$ at the $90\%$ level. Considering the prior is $0.16^{+0.15}_{-0.13}$, this presents a strong case for first-order phase transitions above NS densities. This evidence becomes stronger when the second peak is required (shaded). Furthermore, it is interesting to note that among the subset of samples disfavoring low $\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD}}\right\}$, the preference for the two-peak scenario grows because $\Delta P_{\mathrm{max}}(C_{s,\mathrm{min}}^{\mathrm{UD*}},C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3)$ decreases with increasing $\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD}}\right\}$. The strengthened evidence in the absence of phase transitions is listed in appendix C. Finally, the onset density [66] of the second peak is inferred to lie in the range $n_{B}\in[4.5_{-0.3}^{+0.5},24^{+3}_{-6}]n_{0}$ or $\mu_{B}\in[1.8^{+0.1}_{-0.2},1.9^{+0.2}_{-0.1}]$ GeV when $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, and $n_{B}\in[5.8_{-0.9}^{+1.0},26^{+4}_{-9}]n_{0}$ or $\mu_{B}\in[1.7^{+0.2}_{-0.2},2.0^{+0.3}_{-0.3}]$ GeV if $M_{\text{TOV}}\lesssim 2.2M_{\odot}$. Detailed examination of the magnitude and location of extrema of $C_{s}$ will be reported in a subsequent paper.
@@ IV Discussion
The second peak above NS densities arises because of asymptotic freedom of QCD and indications that EOS in the NS core is stiff. Since the weakly-coupled CQM is almost conformal  [^3] One may not expect large sensitivities towards $X$ considering $C_{s}^{\text{pQCD}}\simeq 1/3$ regardless of its value. But the pressure $P(\mu_{B})$ is an integrated quantity (eqs. 8 and 4) and tiny deviations in $C_{s}$ originated at asymptotic densities cumulate to a considerable spread down to $\mu_{B}\lesssim 3$ GeV. , a featureless and mostly constant $C_{s}=C_{s,\mathrm{mean}}^{\mathrm{UD}}\simeq 1/3$ above NS densities is only feasible if the EOS is also conformal at the center of most massive NSs. Astrophysical observations suggest rapid stiffening of the EOS above $\sim 2-5n_{0}$ (e.g., [39, 7]), yielding relatively low pressure at $\mu_{B}\simeq 1.5-1.9$ GeV. Taken together with a high $P_{\text{pQCD}}$, a minimum $C_{s,\mathrm{min}}^{\mathrm{UD}}<1/3$ preceding a maximum $C_{s,\mathrm{max}}^{\mathrm{UD}}>1/3$ is thus required to bridge a large $\Delta P=P_{\mathrm{CQM}}-P_{\text{TOV}}$. These implications of stiff NS cores are captured by eqs. 10 and 12 rooted in the continuity of $P(\mu_{B})$. A related measure is the trace anomaly $P/\mathcal{E}-1/3$ whose density evolution informs structures in $C_{s}$  [90]. Note though successful matching of the $P(\mathcal{E})$ relation is not always consistent with thermodynamics as $\mu_{B}$ could be discontinuous.
Non-perturbative effects in CQM are strongly indicated by the second peak near the weak-coupling regime because perturbative calculations generally reveal $C_{s}^{\mathrm{CQM}}<1/3$ [50, 51, 52, 53, 54, 55, 56, 57, 58, 55]. Color superconductivity is perhaps the most anticipated mechanism. Since pairing contributions to the sound speed
@@
where $\Delta=\Delta(\mu_{B})$ is the superconducting gap, a maximum exceeding $1/3$ on top of $C_{s}^{\text{pQCD}}\simeq 0.31-0.32$ suggests sizable gaps. While estimates based on both leading-order pQCD and phenomenological models yield the postulated density dependence for $C_{s}^{\mathrm{CQM}}=C_{s}^{\text{pQCD}}+C_{s}^{\Delta}$ since $\Delta\sim\exp(-1/g_{s})$ falls off faster than any powers of $g_{s}$ in $C_{s}^{\text{pQCD}}=1/3+\mathcal{O}(g_{s}^{2})$ as $\mu_{B}\rightarrow\infty$, only the latter predicts sufficiently large gaps $\Delta\gtrsim 100~{}\text{MeV}$ at $\mu_{B}\simeq 1.6-2.4~{}\text{GeV}$. Therefore, if superconducting gap is the underlying mechanism, it likely receives sizable instanton enhancements [91, 18] or is underestimated by leading-order pQCD calculations in the not-so-weakly-coupled regime. Alternatively, the second peak may be attributed to meson condensates [92, 93]. These possibilities will be investigated in detail later.

### v4

@@ Abstract
I show that in addition to the well-known peak inside massive neutron stars, the sound speed in cold dense QCD matter likely exhibits another peak above neutron star densities before it asymptotes to $c_{s}=\sqrt{C_{s}}=\sqrt{1/3}$. Based on the framework reported in arxiv:2408.16738, this approach does not rely on any assumption about the ultra-dense matter not realized in nature. Current multimessenger observation of neutron stars favors the two-peak scenario with a Bayes factor $5.1_{-0.7}^{+0.9}$, where the uncertainties are systematics due to models of neutron star inner cores. This evidence grows to $27_{-8}^{+11}$ if the $2.6M_{\odot}$ component of GW190814 is a neutron star. The trough in $C_{s}$ separating the two peaks is inferred to be below $0.05^{+0.04}_{-0.02}$ (at the $50\%$ level) if $2.6M_{\odot}$ neutron stars exist. The second peak above $1/3$ beyond baryon chemical potential $\mu_{B}=1.6-1.9$ GeV most likely signals non-perturbative effects in cold quark matter, for instance color superconductivity.
@@ I Introduction
The QCD phase diagram at low temperatures and high baryon density remains elusive because the theoretical and experimental tools needed remain rudimentary. Lattice QCD, the only reliable method to calculate non-perturbative phenomena in QCD, is inapplicable as it suffers from the well-known Fermion sign problem  [1, 2, 3] and heavy-ion collisions cannot simultaneously access low temperatures and high densities. Neutron stars (NSs) provide a unique laboratory to study high-density matter at low-temperature and observation of NS mass, radius, and tidal deformability have provided useful constraints on the equation of state (EOS) at baryon number densities $n_{B}\simeq 2-10n_{0}$ where $n_{0}=0.16~{}\text{fm}^{-3}$ (e.g.,[4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]). Above NS densities reliable descriptions in the range $n_{B}\simeq 10-30n_{0}$ are lacking as the strongly interacting cold quark matter (QCM) renders perturbative calculations untractable. Non-perturbative phenomenon in particular color superconductivity [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29] is believed to be active in this regime, although estimates of its strength vary [30, 31, 32, 33, 34, 35].
@@
I now derive a set of stronger conditions that, at the expense of becoming more conservative, are entirely expressible in terms of the EOS at NS and pQCD densities. To begin with, bounds on the magnitude of $C_{s,\mathrm{min}}^{\mathrm{UD}}$ and $C_{s,\mathrm{max}}^{\mathrm{UD}}$ are reported in the preceding letter [67]. A central ingredient is the mean value EOS, the constant sound speed model that directly connects TOV and pQCD points in the baryon number density $n_{B}$ versus baryon chemical potential $\mu_{B}$ plane. It is shown as the dashed black line in fig. 1. Since slopes of $n_{B}(\mu_{B})$ relations are related to the sound speed
@@
To simplify the discussion, in the main text I assume the extrema are unique, i.e. $C_{s,\mathrm{min}}^{\mathrm{UD*}}=C_{s,\mathrm{min}}^{\mathrm{UD}}$ and $C_{s,\mathrm{max}}^{\mathrm{UD*}}=C_{s,\mathrm{max}}^{\mathrm{UD}}$. A proof of the central results relaxing this assumption is presented in  appendix A. The mean value bound eq. 6 can be used to rewrite eq. 2 as
@@
This is a stronger requirement for NS EOSs since $C_{s,\mathrm{mean}}^{\mathrm{UD}}\geq C_{s,\mathrm{min}}^{\mathrm{UD}}$.
@@
For a given NS model, the requirements eqs. 10 and 3 can be simplified depending on the value of $C_{s,\mathrm{mean}}^{\mathrm{UD}}$. When $C_{s,\mathrm{mean}}^{\mathrm{UD}}>1/3$, the mean value bound implies eq. 3, so eqs. 10 and 7 alone suffice in this scenario. I prove in appendix A that this deduction holds even if $C_{s,\mathrm{min}}^{\mathrm{UD}}\neq C_{s,\mathrm{min}}^{\mathrm{UD*}}$ in the presence of multiple extrema.
@@
when $C_{s,\mathrm{mean}}^{\mathrm{UD}}<1/3$. A proof is presented in appendix A. Intuitively, eq. 12 assures a second peak because EOSs obeying it are soft in the ultra-dense regime so must have segments resembling the maximally soft model characterized by a low trough preceding a high peak.
@@ III Bayesian evidence
To assess the evidence for eq. 13 I adopt a Bayesian approach where ab-initio nuclear theories are taken as prior knowledge and astrophysical observations are imposed as constraints. The outer layers of NSs are constructed following the prescription in refs [84, 65]. It centers on ${\chi\text{EFT}}$ calculations of the pure neutron matter up to N3LO  [85, 86] along with $2\sigma$ truncation errors and is informed by empirical constraints on the symmetric nuclear matter. NS inner cores where the density exceeds $n_{\chi\text{EFT}}=2n_{0}$ are modeled by a class of agnostic parameterizations detailed in appendix B. As for the CQM EOS in the perturbative regime, I take N2LO pQCD predictions [52, 51] at $\mu_{B}=2.4$ GeV supplemented by the soft contribution $\sim\mathcal{O}(\alpha_{s}^{3}\log^{2}\alpha_{s})$ at N3LO [57, 58].
@@
in the absence of superconducting gaps [65, 66]. Above, $\Delta P_{\mathrm{min}}$ is based on the maximally stiff model and its expression is given in [67]. Equation 14 is the result of causality and thermodynamic stability, and ensures every sample is physical. To address the phase space volume, I impose equally probable distributions on the criterion $\Delta P-\Delta P_{\mathrm{mean,max}}$ that are flat on each side ($+$ and --), yielding neutral priors without preferences. Other choices of priors are discussed in appendix C.
Table 1 shows Bayes evidence for scenarios where two peaks are required. It is sensitive to the pQCD renormalization scale $X=\bar{\Lambda}/(\mu_{B}/3)$ and reveals diametric predilections in the extreme cases $X=1,4$. Larger $X$’s lend stronger support because they predict higher $P_{\text{pQCD}}$ thus higher $\Delta P$. Marginalization over this uncertainty is achieved by sampling X from a uniform distribution over the commonly assumed range $X\in[1,4]$ and more restrictive priors motivated by the fact that $P_{\text{pQCD}}$ near $X=1$ is susceptible to pairing contributions [65], by resummed leading and next-to-leading soft contributions [56, 60], and by hints from lattice calculations at finite isospin [87, 34, 35]. The results generally favor the two-peak scenario, and are predominantly determined by the NS maximum mass $M_{\text{TOV}}$ where Bayes factors grow rapidly with increasing $M_{\text{TOV}}$. Constraints on NS radii have very limited effects because current uncertainties are large, and NS radii may not robustly inform the EOS near maximum-mass except in extreme scenarios [65]. Gravitational wave constraints on tidal deformability $\Lambda_{1.4M_{\odot}}\lesssim 600-800$ [6, 7, 9, 10, 11] are implied by the N3LO ${\chi\text{EFT}}$ EOS up to $2n_{0}$ [50, 65]. Results employing ${\chi\text{EFT}}$ to lower densities are reported in appendix C.
@@
This framework also allows one to infer the height of the second peak and the preceding trough. Here, in addition to the physical requirement eq. 14, I impose a similar constraint where $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ is sampled randomly from a uniform distribution on $[0,1]$. This injected likelihood excludes NS EOSs demanding $C_{s}^{\mathrm{UD}}>C_{s,\mathrm{max}}^{\mathrm{UD*}}$, so the resulting posterior on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ provides an estimate for the upper bound on the maximum sound speed  [^2] To be clear, $C_{s}^{\mathrm{UD}}$ may surpass this bound but is not required to.. An upper limit on $C_{s,\mathrm{min}}^{\mathrm{UD*}}$ is then obtained for each NS model by solving $\Delta P=\Delta P_{\mathrm{max}}({\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD*}}\right\}},C_{s,\mathrm{max}}^{\mathrm{UD*}})$ when $\Delta P>\Delta P_{\mathrm{mean}}$ and $\Delta P=\Delta P_{\mathrm{min}}({\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD*}}\right\}},C_{s,\mathrm{max}}^{\mathrm{UD*}})$ otherwise, see ref [67].
Figure 2 shows constraints on the magnitude of the peak and the trough. While posteriors on $C_{s,\mathrm{max}}^{\mathrm{UD*}}$ only marginally deviate from the prior suggesting limited constraints on this upper bound, the minimum sound speed is informed. For $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, one finds ${\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD*}}\right\}}=0.05^{+0.08}_{-0.04}$ at the $90\%$ level. Considering the prior is $0.16^{+0.15}_{-0.13}$, this presents a strong case for first-order phase transitions above NS densities. This evidence becomes stronger when the second peak is required (shaded). Furthermore, it is interesting to note that among the subset of samples disfavoring low $\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD}}\right\}$, the preference for the two-peak scenario grows because $\Delta P_{\mathrm{max}}(C_{s,\mathrm{min}}^{\mathrm{UD*}},C_{s,\mathrm{max}}^{\mathrm{UD*}}=1/3)$ decreases with increasing $\mathrm{max}\left\{C_{s,\mathrm{min}}^{\mathrm{UD}}\right\}$. The strengthened evidence in the absence of phase transitions is discussed in appendix C. Finally, the onset density [67] of the second peak is inferred to lie in the range $n_{B}\in[4.5_{-0.3}^{+0.5},24^{+3}_{-6}]n_{0}$ or $\mu_{B}\in[1.8^{+0.1}_{-0.2},1.9^{+0.2}_{-0.1}]$ GeV when $M_{\text{TOV}}\gtrsim 2.6M_{\odot}$, and $n_{B}\in[5.8_{-0.9}^{+1.0},26^{+4}_{-9}]n_{0}$ or $\mu_{B}\in[1.7^{+0.2}_{-0.2},2.0^{+0.3}_{-0.3}]$ GeV if $M_{\text{TOV}}\lesssim 2.2M_{\odot}$. Detailed examination of the magnitude and location of extrema of $C_{s}$ will be reported in a subsequent paper.
@@ IV Discussion
The second peak above NS densities arises because of asymptotic freedom of QCD and indications that EOS in the NS core is stiff  [^3] Earlier work (e.g. [63, 64, 35]) often finds eq. 14 favoring soft NS cores. I clarified model-independently in [65, 67] that softening is only required in the ultra-dense phase. For a discussion of eq. 14 in the context of Bayesian analyses see appendix C. . Since the weakly-coupled CQM is almost conformal  [^4] One may not expect large sensitivities towards $X$ considering $C_{s}^{\text{pQCD}}\simeq 1/3$ regardless of its value. But the pressure $P(\mu_{B})$ is an integrated quantity (eqs. 8 and 4) and tiny deviations in $C_{s}$ originated at asymptotic densities cumulate to a considerable spread down to $\mu_{B}\lesssim 3$ GeV. , a featureless and mostly constant $C_{s}=C_{s,\mathrm{mean}}^{\mathrm{UD}}\simeq 1/3$ above NS densities is only feasible if the EOS is also conformal at the center of most massive NSs. Astrophysical observations suggest rapid stiffening of the EOS above $\sim 2-5n_{0}$ (e.g., [40, 8]), yielding relatively low pressure at $\mu_{B}\simeq 1.5-1.9$ GeV. Taken together with a high $P_{\text{pQCD}}$, a minimum $C_{s,\mathrm{min}}^{\mathrm{UD}}<1/3$ preceding a maximum $C_{s,\mathrm{max}}^{\mathrm{UD}}>1/3$ is thus required to bridge a large $\Delta P=P_{\mathrm{CQM}}-P_{\text{TOV}}$. A related measure is the trace anomaly $P/\mathcal{E}-1/3$ whose density evolution informs structures in $C_{s}$  [91]. Note though successful matching of the $P(\mathcal{E})$ relation is not always consistent with thermodynamics as $\mu_{B}$ could be discontinuous.
Non-perturbative effects in CQM are strongly indicated by the second peak near the weak-coupling regime because perturbative calculations generally reveal $C_{s}^{\mathrm{CQM}}<1/3$ [51, 52, 53, 54, 55, 56, 57, 58, 59, 56]. Color superconductivity is perhaps the most anticipated mechanism. Since pairing contributions to the sound speed is
@@
where $\Delta=\Delta(\mu_{B})$ is the superconducting gap, a maximum exceeding $1/3$ on top of $C_{s}^{\text{pQCD}}\simeq 0.31-0.32$ suggests sizable gaps. While estimates based on both leading-order pQCD and phenomenological models agree with the postulated density dependence of $C_{s}^{\mathrm{CQM}}=C_{s}^{\text{pQCD}}+C_{s}^{\Delta}$ since $\Delta\sim\exp(-1/g_{s})$ falls off faster than any powers of $g_{s}$ in $C_{s}^{\text{pQCD}}=1/3+\mathcal{O}(g_{s}^{2})$ as $\mu_{B}\rightarrow\infty$, only the latter predicts sufficiently large gaps $\Delta\gtrsim 100~{}\text{MeV}$ at $\mu_{B}\simeq 1.6-2.4~{}\text{GeV}$. Therefore, if superconducting gap is the underlying mechanism, it likely receives sizable instanton enhancements [92, 19] or is underestimated by leading-order pQCD calculations in the not-so-weakly-coupled regime. Alternatively, the second peak may be attributed to meson condensates [93, 94]. These possibilities will be investigated in detail later.